what is the value of y in 11y+3=4y+19 and how did you get that answer?
Answer by uma(370)

11y + 3 = 4y + 19
Adding -4y to either side we get,
11y + 3 - 4y = 4y + 19 - 4y
==> 7y + 3 = 19
Adding -3 to either side we have,
7y + 3 - 3 = 19 - 3
==> 7y = 16
==> 7y/7 = 16/7
==> y = 16/7
==> y = 2 2/7

Good Luck!!!